# CrossLMM: Decoupling Long Video Sequences from LMMs via Dual Cross-Attention Mechanisms

### v1

@@ Abstract
The advent of Large Multimodal Models (LMMs) has significantly enhanced Large Language Models (LLMs) to process and interpret diverse data modalities (e.g., image and video). However, as input complexity increases, particularly with long video sequences, the number of required tokens has grown significantly, leading to quadratically computational costs. This has made the efficient compression of video tokens in LMMs, while maintaining performance integrity, a pressing research challenge. In this paper, we introduce CrossLMM, decoupling long video sequences from LMMs via a dual cross-attention mechanism, which substantially reduces visual token quantity with minimal performance degradation. Specifically, we first implement a significant token reduction from pretrained visual encoders through a pooling methodology. Then, within LLM layers, we employ a visual-to-visual cross-attention mechanism, wherein the pooled visual tokens function as queries against the original visual token set. This module enables more efficient token utilization while retaining fine-grained informational fidelity. In addition, we introduce a text-to-visual cross-attention mechanism, for which the text tokens are enhanced through interaction with the original visual tokens, enriching the visual comprehension of the text tokens. Comprehensive empirical evaluation demonstrates that our approach achieves comparable or superior performance across diverse video-based LMM benchmarks, despite utilizing substantially fewer computational resources. Code is available at: https://github.com/shilinyan99/CrossLMM.
@@ 1 Introduction
Large Multimodal Models (LMMs) [42, 41, 44, 31, 26, 38, 74, 18] enhance Large Language Models (LLMs) [1, 54, 66, 15] with visual perception capabilities, demonstrating remarkable proficiency in image-language [24, 25, 77, 12, 17, 23, 16] and video-language [82, 13, 60, 43, 20, 4, 50, 4] tasks. Contemporary research on LMMs predominantly employs an intermediate module, i.e., the projector, to map visual token representations into LLM embedding spaces, which subsequently serve as prefix content for textual tokens into the LLMs as demonstrated in Figure 1 (a). Nevertheless, the increasing complexity of input modalities, particularly long video sequences, generates substantial quantities of visual tokens. This proliferation of tokens necessitates significant computational resources, thereby constraining the applicability of such models in resource-limited environments and long-context scenarios. Consequently, the development of efficient visual compression methodologies that minimize performance degradation has emerged as a critical research imperative, garnering substantial attention from both academic and industrial communities.
In recent literature, a diverse array of methodologies for visual token compression has emerged, which can be categorized into two principal paradigms as follows. 1) Token Merging before LLM [81, 22, 47], as illustrated in Figure 1 (b), predominantly entails the compression of visual tokens through semantic similarity metrics before feeding into LLMs. Nevertheless, this methodology exhibits significant limitations: it fails to adequately identify visual tokens corresponding to relevant textual elements, compromising spatial relationships and attenuating the model’s capacity for comprehensive image interpretation. 2) Token Pruning within LLM [45, 69, 80, 52], as depicted in Figure 1 (c), implements a systematic reduction of visual tokens within the input layers of LLMs. This methodology typically employs the quantification of attention weights between visual and textual tokens, subsequently eliminating those tokens that manifest lower weight values. While this approach effectively illuminates multi-modal interactions, it might instead overlook crucial semantic information within the visual modality. Furthermore, both methods experience significant performance degradation as the number of visual tokens decreases. Consequently, an intriguing question arises: Is it possible to maintain comparable performance while significantly reducing the number of video tokens?
@@
We evaluate CrossLMM on an extensive spectrum of video-based multimodal benchmarks. Our model, utilizing very few tokens (1 or 9 or 16), demonstrates promising performance across various video assessments, including VideoMME [13] and MLVU [82] and so on. These findings substantiate that the implementation of our dual cross-attention module, coupled with a restricted number of visual tokens, enables LMMs to effectively address diverse visual tasks.
@@ 2.1 Large Multimodal Models
Large language models (LLMs) [1, 54, 66, 15] trained on extensive datasets have demonstrated remarkable capabilities in text understanding and generation tasks, establishing the foundation for developing multi-modal LLMs to solve many traditional multi-modal comprehensive tasks [21, 65, 63, 11, 10, 64, 61, 32, 37]. Among large multimodal models (LMMs) [42, 41, 44, 31, 26, 38, 39, 2, 3], LLaVA [34] has emerged as the predominant architecture, valued for its data efficiency and streamlined design. This approach incorporates a projector that effectively bridges visual and language modalities, while employing instruction-tuning for both the projector and the LLM using comprehensive instruction-following datasets. Recent scholarly work [26, 7, 6, 27, 33, 75, 14] has sought to enhance model performance by implementing high-resolution visual inputs. Concurrently, video-based LLMs [78, 79, 35, 62, 46, 8, 48, 51, 30] have advanced through the extension of visual instruction tuning datasets to accommodate video modality. However, the substantial number of tokens generated by high-resolution and video inputs presents significant computational challenges. Consequently, there is a pressing need to develop efficient methods for visual token compression to address these limitations.
@@ 2.2 Visual Token Compression
With high-resolution image and video inputs, the token count has increased substantially, often exceeding textual token quantities by one to two orders of magnitude. Consequently, the efficient compression of visual tokens has emerged as a critical research challenge. Prior approaches [81, 22, 47, 45, 69, 80, 52] addressing this issue can be categorized into two principal directions. The first category encompasses methods that compress tokens generated by visual encoders or implement efficient projections of visual modalities. For instance, LLaMA-VID [31] employs a QFormer [28] architecture to compress visual tokens into a minimal representation of two tokens. Similarly, Deco [68] implements adaptive pooling at the image patch level. However, these approaches lack integration with textual information for guided compression. Alternative methods gradually reduce visual tokens within Large Language Models (LLMs), yet such approaches frequently compromise the semantic integrity of visual information. Therefore, we propose a dual cross-attention module that simultaneously updates text-relevant visual tokens and visually-relevant textual tokens, thereby effectively compressing visual representations while minimizing information loss.
@@ 3.1 Overall Architecture
Our proposed approach, CrossLMM, inspired by current mainstream architecture LLaVA [34], constructs around three core components, including the frame-wise visual encoder, the visual-language projector, and the Large Language Model (LLM), as demonstrated in Figure 2.
Visual Feature Encoding. For visual feature extraction, we employ an image-based encoder rather than video-based architectures, such as VideoMAE [53] or Video-Swin [36]. This approach extracts features from individual frames sequentially. Specifically, we utilize the pre-trained SigLIP [71] vision encoder for encoding visual information. Our methodology is justified by two key considerations: 1) image encoders normally contain better generalization capabilities with larger-scale training, and 2) the temporal information can be retained by concatenating multiple frame representations. Given an input video-text pair, we sample $T$ frames $v\in\mathbb{R}^{T\times 3\times H\times W}$ from the video clip and apply the visual encoder to extract image features. This process can be formulated as:
@@
Initial Token Merge. After that, the bilinear pooling operator $\mathcal{B}:\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N/9\times D}$ is applied preceding the projection operation, performing 3$\times$3 local patch aggregation. This spatially downsamples the token grid from original $27\times 27$ ($N=729$) to $3\times 3$ ($N=9$) resolution through non-overlapping fusion of spatially adjacent patches, while preserving the feature dimension $D$ through parameter-free structural compression. This architecture introduces geometrically meaningful inductive bias through structural token aggregation while maintaining the integrity of inter-patch spatial relations, which also enhances computational efficiency.
Visual-language Projector. The cross-modal projection module $\Phi:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D^{\prime}}$ is architected as a parameterized two-layer MLP ($\text{MLP}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{d_{h}}\rightarrow\mathbb{R}^{D^{\prime}}$), operating on visual tokens $x_{i}\in\mathbb{R}^{D}$ through successive nonlinear transformations (GeLU [19] activation with layer normalization). This learnable transformation achieves bidirectional semantic alignment by establishing: (1) an injective mapping from vision feature space $\mathcal{V}$ to the LLM’s textual embedding manifold $\mathcal{T}$, and (2) a differentiable inverse approximation $\Phi^{-1}_{\text{approx}}$ preserving topological consistency. As the pivotal interface enabling visio-linguistic fusion, $\Phi$ induces latent space isomorphism through $\ell_{2}$-sphere projection regularization, thus constituting the mathematical foundation for constructing joint embedding space.
Figure 3: Implementation Details of Dual Cross-Attention. For a detailed illustration, please refer to Sec. 3.2 and Sec. 3.3.
@@ 3.2 Visual-to-Visual Cross-attention
To boost both the multi-modal and multi-frame feature fusion, we introduce visual-to-visual (V2V) cross attention module for text-condictioned visual aggregation from coarse-grained to fine-grained. For a multimodal vision-language model, the projected and pooled visual feature $\mathcal{V}\in\mathbb{R}^{L_{v}\times D^{\prime}}$ would concat with text feature $\mathcal{T}\in\mathbb{R}^{L_{t}\times D^{\prime}}$ into the LLM, where $D^{\prime}$ denotes the embedding dimension of LLM, and $L_{v}$, $L_{t}$ represents the length of visual and text tokens, respectively. As shown in Figure 3(a), the V2V module take the visual feature $\mathcal{V}$ (pooled and projected), and text features $\mathcal{T}$ as input $\mathcal{I}=[\mathcal{V};\mathcal{T}]$, where [;][;][ ; ] denotes concatenation. we adopt gated cross-modal fusion mechanisms for input tokens to progressively capture multi-modal information. There consists of two stages, self-attention processing and cross-attention fusion. That can be formulated as,
@@
$\mathcal{I}^{\prime}[:L_{v}]=\mathcal{I}^{\prime}[:L_{v}]+\gamma*\text{Attn}(Q,K,V),$ (7)
@@ 3.3 Text-to-Visual Cross-attention
$\mathcal{I}^{\prime}[L_{v}:L_{v}+L_{t}]=\mathcal{I}^{\prime}[L_{v}:L_{v}+L_{t}]+\gamma*\text{Attn}(Q,K,V),$ (10)
@@ 4.1 Experiment settings
Implementation Details. CrossLMM employs SigLIP2 [55] as the vision encoder. Qwen2.5-1.5B-Instruct [66] serves as the large language model (LLM) for CrossLMM-2B, while Qwen2.5-7B-Instruct [66] is utilized for CrossLMM-7B. For the visual-to-visual module, we integrate connections at every $K$ layers within the LLM, following the same approach as the text-to-visual module. During the pretraining phase, we implement a learning rate of $1\times 10^{-4}$, while in the instruction tuning stage, the learning rate is adjusted to $1\times 10^{-5}$, with the exception of the Vision Transformer (ViT) component, which uses $2\times 10^{-6}$. Our model undergoes training only one epoch during both pretraining and instrcution tuning stage. More details can be found in supplementary materails.
Evaluation Benchmarks. To validate CrossLMM’s general capabilities, we evaluate our model on five video understanding benchmarks spanning short video benchmarks, long video benchmarks, and comprehensive benchmarks. These include two short video benchmarks: MVBench [29] and Perception Test [43], and two long video benchmarks: LongVideoBench [60] and MLVU [82], and a comprehensive benchmark, VideoMME [13], covering videos ranging from minute-level to hour-level.

### v2

@@ Abstract
The advent of Large Multimodal Models (LMMs) has significantly enhanced Large Language Models (LLMs) to process and interpret diverse data modalities (e.g., image and video). However, as input complexity increases, particularly with long video sequences, the number of required tokens has grown significantly, leading to quadratically computational costs. This has made the efficient compression of video tokens in LMMs, while maintaining performance integrity, a pressing research challenge. In this paper, we introduce CrossLMM, decoupling long video sequences from LMMs via a dual cross-attention mechanism, which substantially reduces visual token quantity with minimal performance degradation. Specifically, we first implement a significant token reduction from pretrained visual encoders through a simple but effective pooling methodology. Then, within LLM layers, we employ a visual-to-visual cross-attention mechanism, wherein the pooled visual tokens function as queries against the original visual token set. This module enables more efficient token utilization while retaining fine-grained informational fidelity. In addition, we introduce a text-to-visual cross-attention mechanism, for which the text tokens are enhanced through interaction with the original visual tokens, enriching the visual comprehension of the text tokens. Comprehensive empirical evaluation demonstrates that our approach achieves comparable or superior performance across diverse video-based LMM benchmarks, despite utilizing substantially fewer computational resources.
@@ 1 Introduction
Large Multimodal Models (LMMs) [gpt4o, gpt4v, gemini, llamavid, llavaonevision, vlora] enhance Large Language Models (LLMs) [gpt4, llama2, qwen2_5, llama3] with visual perception capabilities, demonstrating remarkable proficiency in image-language [mmecot, mmsearch, mmereal, mme] and video-language [mlvu, videomme, longvideobench, perceptiontest, worldsense] tasks. Contemporary research on LMMs predominantly employs an intermediate module, i.e., the projector, to map visual token representations into LLM embedding spaces, which subsequently serve as prefix content for textual tokens into the LLMs as demonstrated in Figure 1 (a). Nevertheless, the increasing complexity of input modalities, particularly long video sequences, generates substantial quantities of visual tokens. This proliferation of tokens necessitates significant computational resources, thereby constraining the applicability of such models in resource-limited environments and long-context scenarios. Consequently, the development of efficient visual compression methods that minimize performance degradation has emerged as a critical research imperative, garnering substantial attention from both academic and industrial communities.
In recent literature, a diverse array of methodologies for visual token compression has emerged, which can be categorized into two principal paradigms as follows. 1) Token Merging before LLM [aim, huang2024efficient, song2024less], as illustrated in Figure 1 (b), predominantly entails the compression of visual tokens through semantic similarity metrics before feeding into LLMs. Nevertheless, this methodology exhibits significant limitations: it fails to adequately identify visual tokens corresponding to relevant textual elements, compromising spatial relationships and attenuating the model’s capacity for comprehensive image interpretation. 2) Token Pruning within LLM [shang2024llava, ye2024fit, zhao2024accelerating, sun2025lvpruning], as depicted in Figure 1 (c), implements a systematic reduction of visual tokens within the input layers of LLMs. This methodology typically employs the quantification of attention weights between visual and textual tokens, subsequently eliminating those tokens that manifest lower weight values. While this approach effectively illuminates multi-modal interactions, it might instead overlook crucial semantic information within the visual modality. Furthermore, both methods experience significant performance degradation as the number of visual tokens decreases. Consequently, an intriguing question arises: Is it possible to maintain comparable performance while significantly reducing the number of video tokens?
@@
We evaluate CrossLMM on an extensive spectrum of video-based multimodal benchmarks. Our model, utilizing very few tokens (1 or 9 or 16), demonstrates promising performance across various video assessments, including VideoMME [videomme] and MLVU [mlvu] and so on. These findings substantiate that the implementation of our dual cross-attention module, coupled with a restricted number of visual tokens, enables LMMs to effectively address diverse visual tasks.
@@ 2.1 Large Multimodal Models
Large language models (LLMs) [gpt4, llama2, qwen2_5, llama3] trained on extensive datasets have demonstrated remarkable capabilities in text understanding and generation tasks, establishing the foundation for developing multi-modal LLMs. Among large multimodal models (LMMs) [gpt4o, gpt4v, gemini, llamavid, llavaonevision, vlora, ma2024ee, an2024mc, lin2025drawandunderstand], LLaVA [llava] has emerged as the predominant architecture, valued for its data efficiency and streamlined design. This approach incorporates a projector that effectively bridges visual and language modalities, while employing instruction-tuning for both the projector and the LLM using comprehensive instruction-following datasets. Recent scholarly work [llavaonevision, internvl2, internvl2_5, llavainterleave] has sought to enhance model performance by implementing high-resolution visual inputs. Concurrently, video-based LLMs [llavanextvideo, llavavideo, kangaroo, longvila, longvu] have advanced through the extension of visual instruction tuning datasets to accommodate video modality. However, the substantial number of tokens generated by high-resolution and video inputs presents significant computational challenges. Consequently, there is a pressing need to develop efficient methods for visual token compression to address these limitations.
@@ 2.2 Visual Token Compression
With high-resolution image and video inputs, the token count has increased substantially, often exceeding textual token quantities by one to two orders of magnitude. Consequently, the efficient compression of visual tokens has emerged as a critical research challenge. Prior approaches [aim, huang2024efficient, song2024less, shang2024llava, ye2024fit, zhao2024accelerating, sun2025lvpruning] addressing this issue can be categorized into two principal directions. The first category encompasses methods that compress tokens generated by visual encoders or implement efficient projections of visual modalities. For instance, LLaMA-VID [llamavid] employs a QFormer [blip2] architecture to compress visual tokens into a minimal representation of two tokens. Similarly, Deco [deco] implements adaptive pooling at the image patch level. However, these approaches lack integration with textual information for guided compression. Alternative methods gradually reduce visual tokens within Large Language Models (LLMs), yet such approaches frequently compromise the semantic integrity of visual information. Therefore, we propose a dual cross-attention module that simultaneously updates text-relevant visual tokens and visually-relevant textual tokens, thereby effectively compressing visual representations while minimizing information loss.
@@ 3.1 Overall Architecture
Our proposed approach, CrossLMM, inspired by current mainstream architecture LLaVA [llava], constructs around three core components, including the frame-wise visual encoder, the visual-language projector, and the Large Language Model (LLM), as demonstrated in Figure 2.
Visual Feature Encoding.
For visual feature extraction, we employ an image-based encoder rather than video-based architectures, such as VideoMAE [tong2022videomae] or Video-Swin [liu2022videoswin]. This approach extracts features from individual frames sequentially. Specifically, we utilize the pre-trained SigLIP2 [siglip2] vision encoder for encoding visual information. Our methodology is justified by two key considerations: 1) image encoders normally contain better generalization capabilities with larger-scale training, and 2) the temporal information can be retained by concatenating multiple frame representations. Given an input video-text pair, we sample $T$ frames $v\in\mathbb{R}^{T\times 3\times H\times W}$ from the video clip and apply the visual encoder to extract image features. This process can be formulated as:
@@
Initial Token Merge.
After that, the bilinear pooling operator $\mathcal{B}:\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N/9\times D}$ is applied preceding the projection operation, performing $\times3$ local patch aggregation. This spatially downsamples the token grid from original $27\times 27$ ($N=729$) to $3\times 3$ ($N=9$) resolution through non-overlapping fusion of spatially adjacent patches, while preserving the feature dimension $D$ through parameter-free structural compression. This architecture introduces geometrically meaningful inductive bias through structural token aggregation while maintaining the integrity of inter-patch spatial relations, which also enhances computational efficiency.
Visual-language Projector.
The cross-modal projection module $\Phi:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D^{\prime}}$ is architected as a parameterized two-layer MLP ($\text{MLP}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{d_{h}}\rightarrow\mathbb{R}^{D^{\prime}}$), operating on visual tokens $x_{i}\in\mathbb{R}^{D}$ through successive nonlinear transformations (GeLU [hendrycks2016gaussian] activation with layer normalization). This learnable transformation achieves bidirectional semantic alignment by establishing: (1) an injective mapping from vision feature space $\mathcal{V}$ to the LLM’s textual embedding manifold $\mathcal{T}$, and (2) a differentiable inverse approximation $\Phi^{-1}_{\text{approx}}$ preserving topological consistency. As the pivotal interface enabling visio-linguistic fusion, $\Phi$ induces latent space isomorphism through $\ell_{2}$-sphere projection regularization, thus constituting the mathematical foundation for constructing joint embedding space.
@@ 3.2 Visual-to-Visual Cross-attention
To boost both the multi-modal and multi-frame feature fusion, we introduce visual-to-visual (V2V) cross attention module for text-condictioned visual aggregation from coarse-grained to fine-grained. For a multimodal vision-language model, the projected and pooled visual feature $\mathcal{V}\in\mathbb{R}^{L_{v}\times D^{\prime}}$ would concat with text feature $\mathcal{T}\in\mathbb{R}^{L_{t}\times D^{\prime}}$ into the LLM, where $D^{\prime}$ denotes the embedding dimension of LLM, and $L_{v}$, $L_{t}$ represents the length of visual and text tokens, respectively. As shown in Figure 3(a), the V2V module take the visual feature $\mathcal{V}$ (pooled and projected), and text features $\mathcal{T}$ as input $\mathcal{I}=[\mathcal{V};\mathcal{T}]$, where $[;]$ denotes concatenation. we adopt gated cross-modal fusion mechanisms for input tokens to progressively capture multi-modal information. There consists of two stages, self-attention processing and cross-attention fusion. That can be formulated as,
@@
$\mathcal{I}^{\prime\prime}[:L_{v}]=\text{Attn}(Q,K,V)+\gamma*\mathcal{I}^{\prime}[:L_{v}]$ (7)
@@
Figure 3: Implementation Details of Dual Cross-Attention. For a detailed illustration, please refer to Sec. 3.2 and Sec. $3.$
@@ 3.3 Text-to-Visual Cross-attention
$\mathcal{I}^{\prime}[L_{v}:L_{v}+L_{t}]=\text{Attn}(Q,K,V)+\gamma*\mathcal{I}^{\prime}[L_{v}:L_{v}+L_{t}],$ (10)
@@ 4.1 Experiment settings
Table 1: Comprehensive evaluation of video understanding models. Performance comparison across multiple video understanding benchmarks for different categories of models (proprietary, small-size LMMs, general open-source LMMs, and specialized long video LMMs). The CrossLMM model (highlighted) achieves competitive or superior performance while using significantly fewer tokens per frame (1 or 9 or 16). $\dagger$} represents the vision encoder is ViT-G. Bold: Best result. Underline: Second-best result.
Model Size #tokens MVBench PerceptionTest LongVideoBench MLVU VideoMME VideoMME per frame Avg Val Val M-Avg w/o sub. w sub. Avg. Duration 16s 23s 473s 651s 1010s 1010s Proprietary Models GPT4-V [gpt4v] - - 43.7 - 59.1 49.2 59.9 63.3 GPT4-o [gpt4o] - - 64.6 - 66.7 64.6 71.9 77.2 Gemini-1.5-Pro [gemini] - - 60.5 - 64.0 - 75.0 81.3 Small Size LMMs Qwen2-VL [qwen2vl] 2B - 63.2 - - - 55.6 60.4 InternVL2.5 [internvl2_5] 2B 256 68.8 - 46.0 61.4 51.9 54.1 InternVL3.5 [internvl3_5] 2B - 65.9 - 57.4 64.4 58.4 61.9 CrossLMM 2B 1 58.6 58.1 47.3 57.6 55.5 58.9 CrossLMM 2B 9 63.5 63.5 51.8 61.0 59.1 61.3 Open-Source LMMs VideoLLaMA2 [videollama2] 7B 72 54.6 51.4 - 48.5 47.9 50.3 VideoLLaMA2 [videollama2] 72B 72 62.0 57.5 - - 62.4 64.7 VideoChat2-HD [videochat2] 7B 72 62.3 - - 47.9 45.3 55.7 InternVideo2-HD [internvideo2] 7B 72 67.2 63.4 - - 49.4 - IXComposer-2.5 [xcomposer] 7B 400 69.1 34.4 - 37.3 55.8 58.8 InternVL2 [internvl2] 8B 256 65.8 - 54.6 64.0 54.0 56.9 InternVL2 [internvl2] 76B 256 69.6 - 61.1 69.9 61.2 62.8 InternVL2.5 [internvl2_5] 8B 256 72.0 - 60.0 68.9 64.2 66.9 Qwen2-VL [qwen2vl] 7B - 67.0 62.3 - - 63.3 69.0 Qwen2-VL [qwen2vl] 72B - 73.6 68.0 - - 71.2 77.8 Qwen2.5-VL [qwen2_5vl] 7B - 69.6 70.5 56.0 70.2 65.1 71.6 Qwen2.5-VL [qwen2_5vl] 72B - 70.4 73.2 60.7 74.6 73.3 79.1 LLaVA-NeXT-Video [llavanextvideo] 7B 144 53.1 48.8 49.1 - - 46.5 LLaVA-OneVision [llavaonevision] 7B 196 56.7 57.1 56.3 64.7 58.2 61.5 LLaVA-OneVision [llavaonevision] 72B 196 59.4 66.9 61.3 68.0 66.2 69.5 LLaVA-Video [llavavideo] 7B 676 58.6 67.9 58.2 70.8 63.3 69.7 Open-Source Long Video LMMs LLaMA-VID [llamavid] 7B 2 41.9 44.6 - 33.2 25.9 - Kangaroo [kangaroo] 8B 256 61.0 - 54.8 61.0 56.0 57.6 LongVILA [longvila] 7B 196 67.1 58.1 57.1 - 60.1 65.6 LongVA [longva] 7B 144 - - - 56.3 52.6 54.3 LongLLaVA [longllava] 9B 144 49.1 - - - 43.7 - LongVU [longvu] 7B 64 66.9 - - 65.4 - 60.6 CrossLMM 7B 1 62.7 64.8 54.5 64.0 61.3 63.7 CrossLMM 7B 9 68.2 68.4 56.0 67.2 62.6 64.7 CrossLMM  7B 16 68.8 71.4 60.4 70.7 65.4 67.6
Implementation Details. CrossLMM employs SigLIP2 [siglip2] as the vision encoder. Qwen2.5-1.5B-Instruct [qwen2_5] serves as the large language model (LLM) for CrossLMM-2B, while Qwen2.5-7B-Instruct [qwen2_5] is utilized for CrossLMM-7B. For the visual-to-visual module, we integrate connections at every $K$ layers within the LLM, following the same approach as the text-to-visual module. During the pretraining phase, we implement a learning rate of $1\times 10^{-4}$, while in the instruction tuning stage, the learning rate is adjusted to $1\times 10^{-5}$, with the exception of the Vision Transformer (ViT) component, which uses $2\times 10^{-6}$. Our model undergoes training only one epoch during both pretraining and instrcution tuning stage. More details can be found in supplementary materails.
Evaluation Benchmarks. To validate CrossLMM’s general capabilities, we evaluate our model on five video understanding benchmarks spanning short video benchmarks, long video benchmarks, and comprehensive benchmarks. These include two short video benchmarks: MVBench [videochat2] and Perception Test [perceptiontest], and two long video benchmarks: LongVideoBench [longvideobench] and MLVU [mlvu], and a comprehensive benchmark, VideoMME [videomme], covering videos ranging from minute-level to hour-level.